OpenStudy (anonymous):

Convert 55. 8 miles/hour to km/s.

jagr2713 (jagr2713):

you there @cassytaylor92

jagr2713 (jagr2713):

First we need to know that, 1 mile = 1.61 Km And 1 Hour=3600 second So 55.8miles/hour =(55.8*1.61)KM/Hour =(55.8*1.61)/3600 Km/s =0.024955 Km/s Hope this helps when you come back :D

OpenStudy (anonymous):

yeah. thanks for helping me :) @jagr2713

jagr2713 (jagr2713):

Anytime Buddy :D Thanks for choosing Openstudy!